# ThinkSound: Chain-of-Thought Reasoning in Multimodal Large Language Models for Audio Generation and Editing

### v2

@@ 5.1 Experiment Setup
Evaluation Metrics We conduct comprehensive evaluations using both objective and subjective metrics to assess audio quality, text-audio alignment, and video-audio synchronization. For objective evaluation, we compute the Fréchet Distance (FD) (Kilgour et al., 2018; Copet et al., 2024) in the OpenL3 feature space (Cramer et al., 2019; Evans et al., 2024) to measure distribution-level similarity, which we extend to evaluate stereo audio. We also use the Kullback-Leibler (KL) Divergence (Copet et al., 2024) based on predictions from PaSST model ($\text{KL}_{\text{PaSST}}$) and PaNNs ($\text{KL}_{PaNNs}$) to evaluate label-level consistency. Temporal alignment between audio and video is assessed using DeSync predicted by the Synchformer model (Iashin et al., 2024), while semantic alignment with text is measured using the CLAP score (Wu* et al., 2023; Chen et al., 2022), including both $\text{CLAP}_{cap}$ (caption) and $\text{CLAP}_{\text{CoT}}$ (CoT). For subjective evaluation, we collect human ratings using the Mean Opinion Score (MOS) to evaluate perceived audio quality (MOS-Q) and alignment with video and CoT (MOS-A). Further details are provided in the Supplementary Materials C.
Implementation Details For VAE training, we initialize our VAE using the VAE model weights trained on stereo data at 44.1kHz sample rate provided by Stability AI  [^3] https://github.com/Stability-AI/stable-audio-tools. We employ mixed precision training with a batch size of 144 across 24 A800 GPUs for 500,000 steps. Subsequently, following  Evans et al. (2024), we freeze the VAE encoder and train the VAE decoder with a latent mask ratio of 0.1 for an additional 500,000 steps. We use AdamW (Loshchilov & Hutter, 2019) as the optimizer, setting the generator learning rate to 3e-5 and the discriminator learning rate to 6e-5. In the foundation model training phase, we utilize exponential moving average and automatic mixed precision for 100,000 steps on 8 A100 GPUs, with an effective batch size of 256. We adopt a cfg dropout of 0.2 for each modality with a learning rate of 1e-4. During the task-specific fine-tuning stage, we similarly apply exponential moving average and automatic mixed precision for 50,000 steps on 8 A100 GPUs, maintaining an effective batch size of 256. AdamW remains our optimizer of choice, with a learning rate set at 1e-4. We attach the benchmark details of VGGSound, Movie Gen Audio Bench, and AudioCoT test set for stages 2 and 3 into Supplementary Materials A.3.
@@
GT - - - 0.55 0.28 0.45 4.37$\pm$0.21 4.56$\pm$0.19 - -
See&Hear 118.95 2.26 2.30 1.20 0.32 0.35 2.75$\pm$1.08 2.87$\pm$0.99 415M 19.42
V-AURA 46.99 2.23 1.83 0.65 0.23 0.37 3.42$\pm$1.03 3.20$\pm$1.17 695M 14.00
FoleyCrafter 39.15 2.06 1.89 1.21 0.41 0.34 3.08$\pm$1.21 2.63$\pm$0.88 1.20B 3.84
Frieren 74.96 2.55 2.64 1.00 0.37 0.34 3.27$\pm$1.11 2.95$\pm$1.09 159M -
V2A-Mapper 48.10 2.50 2.34 1.23 0.38 0.32 3.31$\pm$1.02 3.16$\pm$1.04 229M -
MMAudio 43.26 1.65 1.40 0.44 0.31 0.40 3.84$\pm$0.89 3.97$\pm$0.82 1.03B 3.01
ThinkSound 34.56 1.52 1.32 0.46 0.33 0.46 4.02$\pm$0.73 4.18$\pm$0.79 1.30B 1.07
w/o CoT Reasoning 39.84 1.59 1.40 0.48 0.29 0.41 3.91$\pm$0.83 4.04$\pm$0.75 1.30B 0.98
@@ 5.2 Main Results
Video-to-Audio Generation Results We compare our foundation model against existing video-to-audio baselines including Seeing and Hearing (Xing et al., 2024), V-AURA (Viertola et al., 2024), FoleyCrafter (Zhang et al., 2024), Frieren (Wang et al., 2024d), V2A-Mapper (Wang et al., 2024a), and MMAudio (Cheng et al., 2024a). From Table 1, we observe three key findings: (1) GT audio exhibits low $\text{CLAP}_{cap}$ scores (0.28), revealing that original VGGSound captions inadequately capture the semantic content and temporal relationships needed for high-fidelity audio generation. (2) ThinkSound outperforms all baselines across most objective metrics and all subjective metrics. Compared to the strongest baseline (MMAudio), our model achieves substantial improvements in audio quality ($\text{KL}_{\text{PaSST}}$: 1.52 vs. 1.65) and semantic alignment ($\text{CLAP}_{\text{CoT}}$: 0.46 vs. 0.40), while maintaining comparable temporal synchronization.
@@
MMAudio 0.45 0.77 3.95$\pm$0.87 3.62$\pm$1.03
MovieGen 0.47 1.00 3.98$\pm$0.77 3.70$\pm$0.96
ThinkSound 0.51 0.76 4.11$\pm$0.74 3.87$\pm$0.82
Subjective evaluations further confirm these improvements (MOS-Q: 4.02 vs. 3.84, MOS-A: 4.18 vs. 3.97). (3) Removing CoT reasoning notably degrades both audio quality and alignment metrics, especially for $\text{CLAP}_{\text{CoT}}$ decreasing from 0.46 to 0.41, confirming that CoT provides crucial information about sound events, their temporal relationships, and acoustic characteristics.
Furthermore, we conduct an out-of-distribution evaluation on the MovieGen Audio Bench (Polyak et al., 2024) to assess the generalization capability of our model. The results in Table 2 show that our ThinkSound model still achieves the best $\text{CLAP}_{\text{CoT}}$ score of 0.51 while DeSync is on par with the best baseline. For subjective evaluation, ThinkSound achieves the best performance in both alignment and fidelity metrics. This demonstrates that ThinkSound exhibits strong generalization capability across different scenarios.
@@
MMAudio 44.46 1.38 0.41 3.61$\pm$0.63 3.64$\pm$0.69
ThinkSound 43.27 1.32 0.48 3.89$\pm$0.52 3.91$\pm$0.53
w/o CoT 45.28 1.34 0.43 3.77$\pm$0.64 3.81$\pm$0.59
Object-Focused Audio Generation and CoT-Guided Audio Editing For object-focused generation, we compare two approaches: (1) MMAudio which does not have the ROI design, and (2) ThinkSound w/o CoT reasoning. Results in Table 3 show that our interactive approach with CoT reasoning achieves significantly better results, demonstrating superior object-specific sound quality and integration with the foundation audio.
@@
AudioLDM-2 61.28 1.94 0.35 3.28$\pm$0.59 3.48$\pm$0.82
DDPM 55.56 1.75 0.39 3.34$\pm$0.28 3.67$\pm$0.56
ThinkSound 34.78 1.45 0.51 3.92$\pm$0.82 3.85$\pm$0.82
w/o CoT 45.78 1.58 0.44 3.53$\pm$0.45 3.52$\pm$0.62
For text-guided audio editing, we compare ThinkSound with AudioLDM-2 (Liu et al., 2024e) and Edit Friendly DDPM (Huberman-Spiegelglas et al., 2024), adapting these models to our experimental setup. As shown in Table 4, ThinkSound outperforms the baselines across all objective and subjective metrics. Specifically, ThinkSound achieves the lowest FD (34.78) and $\text{KL}_{\text{PaSST}}$ (1.45), and the highest CLAP score (0.51), indicating better audio fidelity and semantic relevance. In human evaluation, ThinkSound also gets the best MOS-Q and MOS-A. Removing CoT reasoning clearly drops all metrics, showing the importance of CoT-guided reasoning for text-based audio editing.
@@ 5.3 Ablation Studies
To better understand the contribution of each component in ThinkSound and to validate the effectiveness of our design choices, we conduct comprehensive ablation studies on the VGGSound test set. Mainly focus on: 1) text encoding strategies and 2) multi-modal integration mechanisms. For more ablation and exploratory results, we attach them to Supplementary Materials D.
@@
Text Encoding Strategies. We evaluate different text encoding strategies with or without CoT reasoning and the results are compiled in Table 5. The results show that (1) CoT reasoning substantially improves audio fidelity, with the FD score improving from 39.84 to 37.65 when comparing CLIP-only to T5-based CoT approaches. (2) The integration of contrastive features from CLIP with contextual features from T5 further enhances performance, reducing KL divergence metrics from 1.54 to 1.52 ($\text{KL}_{\text{PaSST}}$) and from 1.35 to 1.32 ($\text{KL}_{\text{PaNNs}}$).
@@
Multi-Modal Integration Mechanisms. We investigate different multi-modal integration mechanisms for video and audio before feeding into the single-stream transformer and the results are displayed in Table 6. We find that (1) the element-wise addition of video and audio features performs better than audio-only input with weak-supervision global conditioning, especially for synchronization metric DeSync decreasing from 0.50 to 0.46. (2) The gated fusion mechanism outperforms simple element-wise addition and audio-only input across all metrics.
@@ 6 Conclusion
This paper presented ThinkSound, a novel CoT reasoning framework for audio generation and editing that decomposed complex audio generation into three interpretable and user-centric stages. Our evaluations showed that ThinkSound outperformed state-of-the-art methods, producing contextually appropriate and temporally precise soundscapes. The framework’s interactive nature enabled users to refine generated audio through intuitive pointing and natural language instructions, addressing the gap between creative intent and automated generation. While challenges remained in capturing nuanced acoustic properties, our AudioCoT dataset and training strategy established a foundation for future research in intelligent audio generation with applications across film, gaming, and social media. Future work we will explore incorporating physical acoustic modeling and developing more sophisticated reasoning capabilities for complex multi-object sound interactions.
@@ A.1 Data Collection and Preprocessing
To ensure high data quality and consistency, we employ a comprehensive preprocessing pipeline. We begin by removing silent audio-video clips to retain only those with meaningful content. For the AudioSet subset (Gemmeke et al., 2017), we further exclude segments containing human voices based on tag information, as our focus is on non-speech audio. All audio-video clips are then segmented into fixed-length intervals of 9.1 seconds, with any shorter clips discarded to maintain uniformity. To achieve a balanced dataset, we maintain an approximately 1:1 ratio between music and sound effect samples, ensuring equal representation of both categories.
@@ A.2 Quality Control for Automated Data Pipeline
For tasks requiring paired audio components, we utilize GPT-4.1-nano to analyze tag categories from VGGSound (Chen et al., 2020) based on two critical criteria. First, we ensure semantic distinctiveness, where tags must be sufficiently distinct to avoid confusion during audio extraction and removal tasks. Second, we verify contextual plausibility, ensuring that the co-occurrence of paired sounds is contextually reasonable within the same acoustic scene. This balanced approach ensures that our audio pairs are both semantically meaningful and practically useful for audio generation tasks.
@@
VGGSound (Chen et al., 2020) Audio-Video Caption 453.6
AudioSet (Gemmeke et al., 2017) Audio-Video Caption 287.5
AudioSet-SL (Hershey et al., 2021) Audio-Text Caption 262.6
Freesound (Fonseca et al., 2017) Audio-Text Caption 1286.6
AudioCaps (Kim et al., 2019) Audio-Text Caption 112.6
@@ A.3 Benchmark Construction
  • Semantic Consistency: The alignment between the audio and the visual content evaluated by Imagebind score (0.3+ for easy, 0.25-0.3 for medium, 0.2-0.25 for hard), and CLAP score between audio and CoT (0.4+ for easy, 0.3-0.4 for medium, 0.2-0.3 for hard).
@@ B.1 Model Configurations
ThinkSound consists of two primary components: a hierarchical variational autoencoder (VAE) for audio compression and reconstruction, and a flow-matching multimodal transformers.
@@
The encoder consists of five convolutional blocks with channel multipliers [1, 2, 4, 8, 16] and strides [2, 4, 4, 8, 8], projecting the stereo waveform into a 128-dimensional latent space. The decoder mirrors this architecture with transposed convolutions to reconstruct 64-dimensional latent representations back into the time-domain waveform.
@@ B.2 Model Architecture
Figure 5: Multi-stream blocks: $F_{v}$ is the video features, $F_{t}$ is the text features, $x_{t}$ is the audio latents, and $c_{g}$ denotes the global condition.
@@ C.1 Objective Metrics
Feature Distribution Alignment: We project both generated and reference audio into the OpenL3 embedding space Cramer et al. (2019); Evans et al. (2024) and compute the Fréchet Distance (FD) Kilgour et al. (2018); Copet et al. (2024) to assess the similarity between their distributional statistics. We chose OpenL3 because it accepts signals of up to 48kHz while VGGish operates at 16kHz, which is more suitable for our 44kHz audio. Following the previous work Evans et al. (2024), we extend the FD to evaluate stereo signals by projecting left and right channels separately and then averaging the results. Moreover, to evaluate whether the generated audio matches the reference in terms of its distribution, we compute the Kullback-Leibler (KL) DivergenceCopet et al. (2024) between class probability distributions predicted by the PaSST modelKoutini et al. (2021) and PaNNs model (Kong et al., 2020) as classifiers.
Temporal Alignment: To evaluate the synchronization between generated audio and its corresponding video, we adopt the DeSync score predicted by the Synchformer model Iashin et al. (2024), following the practice of Cheng et al. (2024a). For each sample, we truncate the video to match the duration of the generated audio and compute the DeSync score using Synchformer, which operates with a 4.8-second context window. Specifically, we extract both the first and last 4.8-second segments from each video-audio pair, calculate the DeSync score for each segment, and report the average as the final temporal alignment metric. Text-Audio Correspondence: To assess the semantic alignment between generated audio and textual prompts, we utilize the CLAP score Wu* et al. (2023); Chen et al. (2022), which measures similarity in a shared audio-text embedding space. Specifically, we report $\text{CLAP}_{\text{cap}}$ for evaluating alignment with the original video captions and $\text{CLAP}_{\text{{CoT}}}$ for alignment with our constructed CoT descriptions. As discussed in Section 5.2, the original VGGSound captions are often low quality and yield lower CLAP scores, whereas our CoT annotations provide richer semantic detail and achieve higher alignment. Consequently, we primarily use the CoT-audio alignment metric in our evaluations, except in Table 1 where caption-based alignment is also reported.

### v3

@@ 5.1 Experiment Setup
Evaluation Metrics We conduct comprehensive evaluations using both objective and subjective metrics to assess audio quality, text-audio alignment, and video-audio synchronization. For objective evaluation, we compute the Fréchet Distance (FD) (kilgour2018fr) in the OpenL3 feature space (cramer2019look; evans2024fast) to measure distribution-level similarity, which we extend to evaluate stereo audio. We also use the Kullback-Leibler (KL) Divergence (copet2024simple) based on predictions from PaSST model ($\text{KL}_{\text{PaSST}}$) and PaNNs ($\text{KL}_{PaNNs}$) to evaluate label-level consistency. Temporal alignment between audio and video is assessed using DeSync predicted by the Synchformer model (iashin2024synchformer), while semantic alignment with text is measured using the CLAP score (laionclap2023; htsatke2022), including both $\text{CLAP}_{cap}$ (caption) and $\text{CLAP}_{\text{CoT}}$ (CoT). For subjective evaluation, we collect human ratings using the Mean Opinion Score (MOS) to evaluate perceived audio quality (MOS-Q) and alignment with video and CoT (MOS-A). Further details are provided in the Appendix C.
Implementation Details For VAE training, we initialize our VAE using the VAE model weights trained on stereo data at 44.1kHz sample rate provided by Stability AI  [^3] https://github.com/Stability-AI/stable-audio-tools. We employ mixed precision training with a batch size of 144 across 24 A800 GPUs for 500,000 steps. Subsequently, following  evans2024fast, we freeze the VAE encoder and train the VAE decoder with a latent mask ratio of 0.1 for an additional 500,000 steps. We use AdamW (DBLP:conf/iclr/LoshchilovH19) as the optimizer, setting the generator learning rate to 3e-5 and the discriminator learning rate to 6e-5. In the foundation model training phase, we utilize an exponential moving average and automatic mixed precision for 100,000 steps on 8 A100 GPUs, with an effective batch size of 256. We adopt a cfg dropout of 0.2 for each modality with a learning rate of 1e-4. During the task-specific fine-tuning stage, we similarly apply exponential moving average and automatic mixed precision for 50,000 steps on 8 A100 GPUs, maintaining an effective batch size of 256. AdamW remains our optimizer of choice, with a learning rate set at 1e-4. We attach the benchmark details of VGGSound, Movie Gen Audio Bench, and AudioCoT test set for stages 2 and 3 into Appendix A.3.
@@
[…]
@@ 5.2 Main Results
Video-to-Audio Generation Results We compare our foundation model against existing video-to-audio baselines including Seeing and Hearing (xing2024seeing), V-AURA (viertola2024temporally), FoleyCrafter (zhang2024foleycrafter), Frieren (wangfrieren), V2A-Mapper (wang2024v2a), and MMAudio (cheng2024taming). From Table 1, we observe three key findings: (1) GT audio exhibits low $\text{CLAP}_{cap}$ scores (0.28), revealing that original VGGSound captions inadequately capture the semantic content and temporal relationships needed for high-fidelity audio generation. (2) ThinkSound outperforms all baselines across most objective metrics and all subjective metrics. Compared to the strongest baseline (MMAudio), our model achieves substantial improvements in audio quality ($\text{KL}_{\text{PaSST}}$: 1.52 vs. 1.65) and semantic alignment ($\text{CLAP}_{\text{CoT}}$: 0.46 vs. 0.40), while maintaining comparable temporal synchronization.
@@
[…]
Subjective evaluations further confirm these improvements (MOS-Q: 4.02 vs. 3.84, MOS-A: 4.18 vs. 3.97). (3) Removing CoT reasoning notably degrades both audio quality and alignment metrics, especially for $\text{CLAP}_{\text{CoT}}$, decreasing from 0.46 to 0.41, confirming that CoT provides crucial information about sound events, their temporal relationships, and acoustic characteristics.
Furthermore, we conduct an out-of-distribution evaluation on the MovieGen Audio Bench (polyak2024movie) to assess the generalization capability of our model. The results in Table 2 show that our ThinkSound model still achieves the best $\text{CLAP}_{\text{CoT}}$ score of 0.51 while DeSync is on par with the best baseline. For subjective evaluation, ThinkSound achieves the best performance in both alignment and fidelity metrics. This demonstrates that ThinkSound exhibits strong generalization capability across different scenarios.
@@
[…]
Object-Focused Audio Generation and CoT-Guided Audio Editing For object-focused generation, we compare two approaches: (1) MMAudio, which does not have the ROI design, and (2) ThinkSound w/o CoT reasoning. Results in Table 3 show that our interactive approach with CoT reasoning achieves significantly better results, demonstrating superior object-specific sound quality and integration with the foundation audio.
@@
[…]
For text-guided audio editing, we compare ThinkSound with AudioLDM-2 (liu2024audioldm) and Edit Friendly DDPM (huberman2024edit), adapting these models to our experimental setup. As shown in Table 4, ThinkSound outperforms the baselines across all objective and subjective metrics. Specifically, ThinkSound achieves the lowest FD (34.78) and $\text{KL}_{\text{PaSST}}$ (1.45), and the highest CLAP score (0.51), indicating better audio fidelity and semantic relevance. In human evaluation, ThinkSound also gets the best MOS-Q and MOS-A. Removing CoT reasoning clearly drops all metrics, showing the importance of CoT-guided reasoning for text-based audio editing.
@@ 5.3 Ablation Studies
To better understand the contribution of each component in ThinkSound and to validate the effectiveness of our design choices, we conduct comprehensive ablation studies on the VGGSound test set. Mainly focus on: 1) text encoding strategies and 2) multi-modal integration mechanisms. For more ablation and exploratory results, we attach them to Appendix D.
@@
Text Encoding Strategies. We evaluate different text encoding strategies with or without CoT reasoning, and the results are compiled in Table 5. The results show that (1) CoT reasoning substantially improves audio fidelity, with the FD score improving from 39.84 to 37.65 when comparing CLIP-only to T5-based CoT approaches. (2) The integration of contrastive features from CLIP with contextual features from T5 further enhances performance, reducing KL divergence metrics from 1.54 to 1.52 ($\text{KL}_{\text{PaSST}}$) and from 1.35 to 1.32 ($\text{KL}_{\text{PaNNs}}$).
@@
Multi-Modal Integration Mechanisms. We investigate different multi-modal integration mechanisms for video and audio before feeding them into the single-stream transformer, and the results are displayed in Table 6. We find that (1) the element-wise addition of video and audio features performs better than audio-only input with weak-supervision global conditioning, especially for synchronization metric DeSync, decreasing from 0.50 to 0.46. (2) The gated fusion mechanism outperforms simple element-wise addition and audio-only input across all metrics.
@@ 6 Conclusion
This paper presented ThinkSound, a novel CoT reasoning framework for audio generation and editing that decomposed complex audio generation into three interpretable and user-centric stages. Our evaluations showed that ThinkSound outperformed state-of-the-art methods, producing contextually appropriate and temporally precise soundscapes. The framework’s interactive nature enabled users to refine generated audio through intuitive pointing and natural language instructions, addressing the gap between creative intent and automated generation. While challenges remained in capturing nuanced acoustic properties, our AudioCoT dataset and training strategy established a foundation for future research in intelligent audio generation with applications across film, gaming, and social media. In future work, we will explore incorporating physical acoustic modeling and developing more sophisticated reasoning capabilities for complex multi-object sound interactions.
Acknowledgements
The research was supported by the NSFC (No. 62206234) of Mainland China.
@@ A.1 Data Collection and Preprocessing
To ensure high data quality and consistency, we employ a comprehensive preprocessing pipeline. We begin by removing silent audio-video clips to retain only those with meaningful content. For the AudioSet subset (gemmeke2017audio), we further exclude segments containing human voices based on tag information, as our focus is on non-speech audio. All audio-video clips are then segmented into fixed-length intervals of 9.1 seconds, with any shorter clips discarded to maintain uniformity. To achieve a balanced dataset, we maintain an approximately 1:1 ratio between music and sound effect samples, ensuring equal representation of both categories.
@@ A.2 Quality Control for Automated Data Pipeline
For tasks requiring paired audio components, we utilize GPT-4.1-nano to analyze tag categories from VGGSound (chen2020vgg) based on two critical criteria. First, we ensure semantic distinctiveness, where tags must be sufficiently distinct to avoid confusion during audio extraction and removal tasks. Second, we verify contextual plausibility, ensuring that the co-occurrence of paired sounds is contextually reasonable within the same acoustic scene. This balanced approach ensures that our audio pairs are both semantically meaningful and practically useful for audio generation tasks.
@@
VGGSound (chen2020vgg) Audio-Video Caption 453.6
AudioSet (gemmeke2017audio) Audio-Video Caption 287.5
AudioSet-SL (hershey2021benefit) Audio-Text Caption 262.6
Freesound (fonseca2017freesound) Audio-Text Caption 1286.6
AudioCaps (kim2019audiocaps) Audio-Text Caption 112.6
@@ A.3 Benchmark Construction
  • Semantic Consistency: The alignment between the audio and the visual content is evaluated by the ImageBind score (0.3+ for easy, 0.25-0.3 for medium, 0.2-0.25 for hard), and the CLAP score between audio and CoT (0.4+ for easy, 0.3-0.4 for medium, 0.2-0.3 for hard).
@@ B.1 Model Configurations
ThinkSound consists of two primary components: a hierarchical variational autoencoder (VAE) for audio compression and reconstruction, and a flow-matching multimodal transformer.
@@
The encoder consists of five convolutional blocks with channel multipliers [1, 2, 4, 8, 16] and strides [2, 4, 4, 8, 8], projecting the stereo waveform into a 128‐dimensional latent space. The decoder mirrors this architecture with transposed convolutions to reconstruct 64‐dimensional latent representations back into the time‐domain waveform.
@@ B.2 Model Architecture
Figure 5: Multi‐stream blocks: $F_{v}$ is the video features, $F_{t}$ is the text features, $x_{t}$ is the audio latents, and $c_{g}$ denotes the global condition.
@@ C.1 Objective Metrics
Feature Distribution Alignment: We project both generated and reference audio into the OpenL3 embedding space cramer2019look; evans2024fast and compute the Fréchet Distance (FD) kilgour2018fr; copet2024simple to assess the similarity between their distributional statistics. We chose OpenL3 because it accepts signals of up to 48kHz while VGGish operates at 16kHz, which is more suitable for our 44kHz audio. Following the previous work evans2024fast, we extend the FD to evaluate stereo signals by projecting left and right channels separately and then averaging the results. Moreover, to evaluate whether the generated audio matches the reference in terms of its distribution, we compute the Kullback-Leibler (KL) Divergencecopet2024simple between class probability distributions predicted by the PaSST modelkoutini2021efficient and PaNNs model (kong2020panns) as classifiers.
Temporal Alignment: To evaluate the synchronization between generated audio and its corresponding video, we adopt the DeSync score predicted by the Synchformer model iashin2024synchformer, following the practice of cheng2024taming. For each sample, we truncate the video to match the duration of the generated audio and compute the DeSync score using Synchformer, which operates with a 4.8-second context window. Specifically, we extract both the first and last 4.8-second segments from each video-audio pair, calculate the DeSync score for each segment, and report the average as the final temporal alignment metric. Text-Audio Correspondence: To assess the semantic alignment between generated audio and textual prompts, we utilize the CLAP score laionclap2023; htsatke2022, which measures similarity in a shared audio-text embedding space. Specifically, we report $\text{CLAP}_{\text{cap}}$ for evaluating alignment with the original video captions and $\text{CLAP}_{\text{{CoT}}}$ for alignment with our constructed CoT descriptions. As discussed in Section 5.2, the original VGGSound captions are often low quality and yield lower CLAP scores, whereas our CoT annotations provide richer semantic detail and achieve higher alignment. Consequently, we primarily use the CoT-audio alignment metric in our evaluations, except in Table 1 where caption-based alignment is also reported.